if the number of dimes are 4 less than 5 times the quarter find how many of each type does he have
and the total is $1.85

Answer by oberobic(2304)

d = number of dimes
10d = value of dimes in cents
q = number of quarters
25q = value of quarters in cents
.
The number of dimes is 4 less than 5 times the number of quarters.
d = 5*q -4
.
10d + 25q = 185 cents
.
Substitute
.
10*(5q-4) + 25q = 185
.
50q -40 +25q = 185
.
75q = 225
.
q = 3
.
d = 5*3-4 = 11
.
Check the values.
25*3 = 75 cents
11*10 = 110 cents
75+110 = 185 cents
Correct.
.
Answer: 3 quarters and 11 dimes.
.
Done.